If x = 1 and x = -8 then form a quadratic equation.: If x = 1 and x = -8 then form a quadratic equation.
Answer by jim_thompson5910(9166)

x=1 or x=-8 Start with the given solutions


x-1=0 or x+8=0 Get all of the terms to the left side to make the right side equal to zero


(x-1)(x+8)=0 Use the zero product property to join the two equations.


x^2+8x-x-8=0 FOIL


x^2+7x-8=0 Combine like terms.


So x^2+7x-8=0 has the solutions x=1 or x=-8